OpenStudy (anonymous):

Medal given to person who answers! Find the area of a sector of a circle with radius 12 and arc length 10(pi). Show work

OpenStudy (anonymous):

Ok for this one is is easiest to draw it out first.

OpenStudy (anonymous):



OpenStudy (anonymous):

That is the final answer?

OpenStudy (anonymous):

No, right?

OpenStudy (anonymous):

Like what are the steps?

OpenStudy (anonymous):

No that isn't the answer.. You have to plug in the information they have given you. The r in this formula represents 12, the radius. While m represents the measure of the length they gave you, 10pi.

OpenStudy (anonymous):

So:\[144 \left( 10\pi \right) \over 360\]

OpenStudy (anonymous):

1440pi over 360?

OpenStudy (anonymous):

a=pir^2*10pi/2pir=5pir

OpenStudy (anonymous):

Could someone just write out the entire answer so that i can understand the steps easier? Thanks

OpenStudy (anonymous):

@Emah Could you help?

OpenStudy (anonymous):

A=pi r^2/2 pi r*10 pi=5 pi r=5 pi 12=60 pi

OpenStudy (anonymous):

Sorry but im not following..Maybe make it easier to read?

OpenStudy (anonymous):

I'm sorry, I know how to do this, but I can't figure out the 10pi being the measure.. Usually its a degree..

OpenStudy (anonymous):

A=(πr^2)/2πr*10π=5πr=5π*12=60π

OpenStudy (anonymous):

does the / mean a fraction like (pi)(r)^2 over something else?